# FutureSightDrive: Thinking Visually with Spatio-Temporal CoT for Autonomous Driving

### v1

@@ 4.2 Main results
We present trajectory planning performance on nuScenes following previous methods [20, 16] in Table 1. When using ego status, FSDrive surpasses previous SOTA methods using ego status in ST-P3 and UniAD metrics. However, following BEV-Planner [25] findings about ego-status’s performance boost, we prioritize non-ego-status evaluations. Compared to non-autoregressive (e.g., UniAD) and autoregressive methods (e.g., OminiDrive), FSDrive demonstrates superior effectiveness. Notably, FSDrive outperforms Doe-1 [72] which also enables vision generation (L2: 0.53 vs. 0.70 and 0.96 vs. 1.26; collision: 0.19 vs. 0.21 and 0.40 vs. 0.53), indicating limitations in their VQ-VAE-based discrete visual features for understanding. For a fair comparison, we also used LLaVA like methods [50, 17, 65]. Under the corresponding settings, FSDrive still has excellent competitiveness, indicating that FSDrive can be widely applied to any existing MLLM.
@@
Although we generate future frames as CoT for trajectory planning, we still validate visual quality via FID in Table 2. To enable rapid generation for real-time driving, we generate frames at 128$\times$192 resolution. Our autoregressive FSDrive achieves competitive performance against specialized diffusion models. Compared to Doe-1 [72] which employs the vision generation MLLM Lumina-mGPT 7B [26], FSDrive 2B maintains superior advantages, indicating that the visual generation capabilities of MLLM can be effectively unlocked even with minimal data.
@@
FSDrive’s scene understanding was evaluated on DriveLM in Table 3, achieving 0.57 and outperforming recent methods like Cube-LLM [5] and OminiDrive [50]. This highlights the effectiveness of FSDrive pre-training paradigm for generation and understanding.
@@
DriveLM baseline [ECCV24] [39] 0.00 0.65 0.05 0.08 0.10 0.28 0.32
Cube-LLM [ICLR25] [5] 0.39 0.89 0.16 0.20 0.31 0.39 0.50
TrackingMeetsLMM [arxiv25] [19] 0.60 0.58 0.72 0.72 0.04 0.36 0.52
SimpleLLM4AD [arxiv24] [70] 0.66 0.57 0.76 0.73 0.15 0.35 0.53
OminiDrive [CVPR25] [50] 0.70 0.65 0.52 0.73 0.13 0.37 0.56
@@ 4.3 Ablation study
We conduct ablation studies on future frames generation in Table 6(b). Table 6(a) shows that larger pre-training datasets improve MLLM’s visual generation capability. Despite being much smaller (200K vs. 100M in previous work [54, 43, 9]), our data achieves more robust visual generation. Scaling datasets may further enhance performance. Table 6(b) confirms our progressive method improves autoregressive image generation.
@@ 5 Conclusion
This paper proposes FSDrive, an autonomous driving framework based on spatio-temporal CoT that enables VLMs to think visually. By unifying future scene generation and perception results through intermediate image-form reasoning steps, our FSDrive eliminates the semantic gap caused by cross-modal conversions and establishes an end-to-end visual reasoning pipeline. The VLM serves dual roles: as a world model that predicts future image frames with lane divider and 3D detection, and as an inverse dynamics model that plans trajectory based on both current observations and future predictions. To enable visual generation in VLMs, we present a pretraining paradigm that unifies visual generation and understanding, along with a progressive easy-to-hard visual CoT to enhance autoregressive image generation. Extensive experimental results demonstrate the effectiveness of the proposed FSDrive method, advancing autonomous driving towards visual reasoning.
@@
Though autonomous driving requires surrounding environmental awareness, considering real-time efficiency, we currently only generate future frames for the front-view. Future work can attempt to generate Surround views to achieve safer autonomous driving. In terms of impact, the ethical challenges posed by LLMs extend to autonomous driving. Advances in technology and regulation will drive development of safer, more efficient systems.

### v2

@@ 4.2 Main results
We present trajectory planning performance on nuScenes following previous methods [29, 21] in Table 1. When using ego status, FSDrive surpasses previous SOTA methods using ego status in ST-P3 and UniAD metrics. However, following BEV-Planner [34] findings about ego-status’s performance boost, we prioritize non-ego-status evaluations. Compared to non-autoregressive (e.g., UniAD) and autoregressive methods (e.g., OmniDrive), FSDrive demonstrates superior effectiveness. Notably, FSDrive outperforms Doe-1 [95] which also enables vision generation (L2: 0.53 vs. 0.70 and 0.96 vs. 1.26; collision: 0.19 vs. 0.21 and 0.40 vs. 0.53), indicating limitations in their VQ-VAE-based discrete visual features for understanding. For a fair comparison, we also used LLaVA like methods [64, 26, 86, 75]. Under the corresponding settings, FSDrive still has excellent competitiveness, indicating that FSDrive can be widely applied to any existing MLLM.
Method DriveGAN DriveDreamer Drive-WM GenAD GEM Doe-1 FSDrive
 [CVPR21 [30]]  [ECCV24 [65]]  [CVPR24 [66]]  [CVPR24 [77]]  [CVPR25 [16]]  [arxiv24 [95]]
[…]
FID $\downarrow$ 73.4 52.6 15.8 15.4 10.5 15.9 10.1
Results on NAVSIM.
Table 2 shows the evaluation results for NAVSIM [10]. All approaches rely exclusively on camera input, with no lidar data being used. Achieving a PDMS score of 85.1, FSDrive outperforms prior camera-only methods like LAW [33] and DiffusionDrive-Cam [36], thus showcasing its efficacy in the pseudo closed-loop setting.
@@
[…]
@@
FSDrive’s scene understanding was evaluated on DriveLM in Table 4, achieving 0.57 and outperforming recent methods like Cube-LLM [7] and OmniDrive [64]. This highlights the effectiveness of FSDrive pre-training paradigm for generation and understanding.
@@
DriveLM baseline [54] 0.00 0.65 0.05 0.08 0.10 0.28 0.32
Cube-LLM [7] 0.39 0.89 0.16 0.20 0.31 0.39 0.50
TrackingMeetsLMM [28] 0.60 0.58 0.72 0.72 0.04 0.36 0.52
SimpleLLM4AD [93] 0.66 0.57 0.76 0.73 0.15 0.35 0.53
OmniDrive [64] 0.70 0.65 0.52 0.73 0.13 0.37 0.56
@@ 4.3 Ablation study
We conduct ablation studies on future frames generation in Table 7(b). Table 7(a) shows that larger pre-training datasets improve MLLM’s visual generation capability. Despite being much smaller (200K vs. 100M in previous work [70]), our data achieves more robust visual generation. Scaling datasets may further enhance performance. Table 7(b) confirms our progressive method improves autoregressive image generation.
@@ 5 Conclusion
This paper proposes FSDrive, an autonomous driving framework based on spatio-temporal CoT that enables VLAs to think visually. By unifying future scene generation and perception results through intermediate image-form reasoning steps, our FSDrive eliminates the semantic gap caused by cross-modal conversions and establishes an end-to-end visual reasoning pipeline. The VLA serves dual roles: as a world model that predicts future image frames with lane divider and 3D detection, and as an inverse dynamics model that plans trajectory based on both current observations and future predictions. To enable visual generation in VLAs, we present a pretraining paradigm that unifies visual generation and understanding, along with a progressive easy-to-hard visual CoT to enhance autoregressive image generation. Extensive experimental results demonstrate the effectiveness of the proposed FSDrive method, advancing autonomous driving towards visual reasoning.
@@
Though autonomous driving requires surrounding environmental awareness, considering real-time efficiency, we currently only generate future frames for the front-view. Future work can attempt to generate Surround views to achieve safer autonomous driving. Moreover, more robust visual quality can be achieved in future work through the use of larger training datasets and a more advanced unified paradigm that integrates generation and understanding. In terms of impact, the ethical challenges posed by LLMs extend to autonomous driving. Advances in technology and regulation will drive development of safer, more efficient systems.
Acknowledgments.
This work was support by the National Natural Science Foundation of China No. 62572385, the Fundamental Research Funds for the Central Universities No. xxj032023020, and CAAI-CANN Open Fund, developed on OpenI Community.